# Object Segmentation-Assisted Inter Prediction for Versatile Video Coding

### v1

@@ Abstract
In modern video coding standards, block-based inter prediction is widely adopted, which brings high compression efficiency. However, in natural videos, there are usually multiple moving objects of arbitrary shapes, resulting in complex motion fields that are difficult to compactly represent. This problem has been tackled by more flexible block partitioning methods in the Versatile Video Coding (VVC) standard, but the more flexible partitions require more overhead bits to signal and still cannot be made arbitrary shaped. To address this limitation, we propose an object segmentation-assisted inter prediction method (SAIP), where objects in the reference frames are segmented by some advanced technologies. With a proper indication, the object segmentation mask is translated from the reference frame to the current frame as the arbitrary-shaped partition of different regions without any extra signal. Using the segmentation mask, motion compensation is separately performed for different regions, achieving higher prediction accuracy. The segmentation mask is further used to code the motion vectors of different regions more efficiently. Moreover, segmentation mask is considered in the joint rate-distortion optimization for motion estimation and partition estimation to derive the motion vector of different regions and partition more accurately. The proposed method is implemented into the VVC reference software, VTM version 12.0. Experimental results show that the proposed method achieves up to 1.98%, 1.14%, 0.79%, and on average 0.82%, 0.49%, 0.37% BD-rate reduction for common test sequences, under the Low-delay P, Low-delay B, and Random Access configurations, respectively.
@@
Inter prediction, motion compensation, motion vector coding, motion estimation, partition estimation, object segmentation, video coding, VVC.I Introduction
@@
Recently, with the repaid development of segmentation technologies[15, 16, 17], moving objects can be pixel-accurately segmented in video frames. Inspired by them, we rethink the assistance of segmentation applied in inter prediction, and sufficiently explore the guidance of accurate object-aware partition information for prediction and coding process in codec. In this paper, we propose an object segmentation-assisted inter prediction (SAIP) method that does not restrict the partitioning shape and further takes advantage of the segmentation information to assist the entire inter prediction process, including MC, motion vector coding (MVC), and ME.
@@ II-A2 Line-based Geometric Partitioning Schemes
Line-based geometric partitioning schemes (GEO) have been intensively studied to capture the non-uniform motion field [10, 18] during the development of HEVC and VVC. In GEO, a predefined straight line is used to split the rectangular block into two wedge-shaped sub-parts. Each sub-part has its MV to describe the motion field and performs MC to obtain the prediction. In VVC, merge-mode-based GEO with 64 partitioning modes has been adopted [11]. The mode is selected for a coding unit from the predefined 64 partition line candidates, and the description information of partition line is represented in the form of the offset parameter and angle parameter . For motion information of sub-parts, only unidirectional motion information is used, and the GEO merge list is derivated from the regular merge list by the parity of the GEO merge index. Although the geometric partition can divide the motion-inconsistent sub-parts in a block to perform MC individually, the straight line can only approximate the discontinuities of the motion field. Meanwhile, bit consumption for partition line representation is unavoidable.
@@ II-A3 Segmentation-based Partitioning Schemes
To represent the motion field of arbitrary-shaped object accurately in natural videos, segmentation-based partitioning schemes are tentatively introduced into existing video coding standards [19, 20]. In the segmentation-based partition, the partition is mainly derived from the reference frame/block by applying a segmentation algorithm at the encoder/decoder.
@@ III Proposed Framework
Figure 2 shows our proposed object segmentation-assisted inter prediction (SAIP) framework from the perspective of decoding. In SAIP, the key lies in that the segmentation information is utilized to assist the overall prediction and coding process in codec. It mainly consists of four modules, including MC (Section IV), MVC (Section V), rate-distortion optimization for ME and partition estimation (Section VI), segmentation for reference pictures (Section III of supplementary material). The difference between the encoder and decoder is only the additional process of ME and partition estimation. In this section, we briefly introduce the process of SAIP in the decoder.
@@
In the decoding process (Fig. 2 ), different from the previous inter prediction methods, the coding block is divided into multiple motion-inconsistent regions (primary region, secondary region) to match the prediction in SAIP. First, the motion data of the current coding block is obtained by parsing the bitstream. Concerning the motion data for the coding block, two MVs (MV, MV) and reference picture indexes (Ref, Ref) correspond to the motion of different regions (primary motion, secondary motion). The partition of the coding block is translated from the segmentation mask of the reference block by the guidance of MV. Second, the segmentation mask is utilized to derive the motion information of the secondary region (MV). Third, with the assistance of the segmentation mask, the final prediction can be derived by MC based on MVs of different regions.
IV Segmentation-Assisted Motion Compensation
@@
In our proposed segmentation-assisted framework, we introduce efficient deep learning-based segmentation methods, SCNet[27] and STCN[33], to generate the segmentation. The reference frame can be segmented with the same setup in the encoder and decoder (the detail for design and implementation of segmentation in codec is mentioned in Section III of supplementary material). The segmentation mask of the reference frame is generated when the coding frame is reconstructed, and stored in buffer memory for segmentation-assisted inter prediction of subsequent frames. With a proper indication, the object segmentation mask of the reference block is translated to the coding block as the arbitrary-shaped partition of motion-inconsistent regions. In this section, we assume that the partition and the motion information of different regions are obtained, and introduce how to utilize the segmentation mask to assist the MC for different motion-inconsistent regions (sub-regions) within the block.
@@
For different sub-regions, two MVs (MV and MV) are used to represent their motion and correspond to two block-level reference blocks named primary prediction and secondary prediction (P, P). Considering the accurate representation of motion information, we propose segmentation-assisted motion compensation (SA-MC) to generate the prediction of different sub-regions within a block, the process of SA-MC is shown in Fig. 3 . In particular, the segmentation mask assists the two steps of MC, including the integer and fractional parts.
IV-A Segmentation-Assisted Integer-pixel Motion Compensation
@@
On both sides of the non-linear partition line, the 2-pixel range region is defined as the edge region. For the edge region to be predicted, not only its own MV but also the MV of the neighboring region can be used to derive the prediction signal. Fig. 4 shows a set of ORMC examples. A blending process is applied to the pixels around the non-linear partition line by a distance-based adaptive weighted strategy. Pixels are classified into three categories: one pixel away from the partition line, two pixels away, and more than two pixels away. Each category corresponds to a different weight for the fusion of P and P in every pixel.
@@
IV-B Segmentation-Assisted Fractional-pixel Motion Compensation
@@
V Segmentation-Assisted Motion Vector Coding
@@
In SAIP, the motion of each coding block is divided into two parts (primary motion and secondary motion) to derive the motion data, which corresponds to two MVs (MV, MV). One is coded by the index of the primary candidate list (length: 70), the list includes regular merge candidates (Merge) and merge with motion vector difference (MMVD) candidates. The other one is coded by the index of the secondary candidate list (length: 7), the list is established with the assistance of segmentation mask.
@@
In the coding process, the primary candidate list is established at first for the coding block, and each candidate can provide a MV (MV) for the current coding block’s rough motion representation. Due to multiple motions in the coding block, MV can only consider describing a part of the motion. To further describe the multiple motions, the partition of motion area is needed. By the guidance of MV, the segmentation mask of the primary reference block can be translated to the current coding block as the partition of the motion area. From the mask, the coding block can be partitioned into two motion-inconsistent regions (primary region, secondary region). For the motion of the primary region, it can be represented by MV. In order to efficiently represent the motion of other regions, we propose a segmentation-assisted candidate derivation (SA-CD) strategy for the MV candidate derivation of other parts.
@@
The construction of the regular merge candidate list follows the fixed predefined order to traverse the potential candidates (B1A1B0A0B2T…) to fill the candidate list. Figure 5 .(i) shows the position of spatial and temporal candidates for the current coding unit (CU), which is used to copy the motion information of the neighboring block for reference. In SA-CD, the segmentation mask (object-aware partition) is utilized as the criterion to perform motion analysis for coding block and further adaptively select and rank the high-probability candidates to construct the candidate list, thereby considering different qualities of the potential MV candidates. The example is shown in Figure 5 .(ii). First, the distribution pattern of the secondary region in the partition of the coding block’s motion-inconsistent regions is searched. The nearest four-pixel values near the corner are taken as the classification rule. In that case, the four pixels in the upper left and upper right corner of the segmentation mask are all zero (black). It is considered that these corners of the segmentation mask are in the secondary region. Similarly, the four corners can be in the secondary or the primary region, so the distribution of the motion-inconsistent regions can be classified into 16 patterns, all patterns are shown in Fig.1 of supplementary material. Different patterns of distribution of motion-inconsistent regions correspond to different candidate lists. The secondary candidate list is also derived from the candidates in Fig. 5 .(i). The construction order is determined by the distribution pattern of secondary region, which assumes that the motion information of neighboring blocks closer to the secondary region is more likely to match the motion information of secondary region. Therefore, according to the distribution pattern (blue dotted box) of Fig. 5 .(ii), the order can be determined as B1B2B0(0,0)A0A1T. The candidate list of all distribution patterns is also shown in Fig.1 of supplementary material.
@@
VI Segmentation-Assisted Rate-Distortion Optimization for Motion Estimation and
Partition Estimation
@@ VI Segmentation-Assisted Rate-Distortion Optimization for Motion Estimation and
For the ME of different motion-inconsistent regions, the accuracy of motion representation depends on whether the partition (segmentation mask) of the motion area can approximate the actual motion field in the coding block. In the previous related work (GEO[11], PCS-16[14], TIP-19[13]), template-based or hypothesis-based partition derivation methods are utilized, the fixed search range of partition varieties and individual RDO process (for ME and partition search) limit the partition search space of coding block and the accuracy of MVs, thereby further impacting on the prediction quality. To solve these limits, segmentation-assisted rate-distortion optimization (SA-RDO) is applied to the proposed SAIP framework. In Fig. 6 , the object segmentation-assisted inter prediction framework with proposed SA-RDO is shown from the perspective of encoding. The motion candidates of MV imply various partition candidates are used to suppose the motion area partition of current coding block (partition information does not need to be explicitly signaled). And the partition estimation and the ME of primary motion and secondary motion are joint to search the best combination in the whole RDO process.

### v2

@@ Abstract
In modern video coding standards, block-based inter prediction is widely adopted, which brings high compression efficiency. However, in natural videos, there are usually multiple moving objects of arbitrary shapes, resulting in complex motion fields that are difficult to represent compactly. This problem has been tackled by more flexible block partitioning methods in the Versatile Video Coding (VVC) standard, but the more flexible partitions require more overhead bits to signal and still cannot be made arbitrarily shaped. To address this limitation, we propose an object segmentation-assisted inter prediction method (SAIP), where objects in the reference frames are segmented by some advanced technologies. With a proper indication, the object segmentation mask is translated from the reference frame to the current frame as the arbitrary-shaped partition of different regions without any extra signal. Using the segmentation mask, motion compensation is separately performed for different regions, achieving higher prediction accuracy. The segmentation mask is further used to code the motion vectors of different regions more efficiently. Moreover, the segmentation mask is considered in the joint rate-distortion optimization for motion estimation and partition estimation to derive the motion vector of different regions and partition more accurately. The proposed method is implemented into the VVC reference software, VTM version 12.0. Experimental results show that the proposed method achieves up to 1.98%, 1.14%, 0.79%, and on average 0.82%, 0.49%, 0.37% BD-rate reduction for common test sequences, under the Low-delay P, Low-delay B, and Random Access configurations, respectively.
@@
Inter prediction, motion compensation, motion estimation, motion vector coding, object segmentation, partition estimation, video coding, VVC.I Introduction
@@
Recently, with the repaid development of segmentation technologies[15, 16, 17, 18], moving objects can be pixel-accurately segmented in video frames. Inspired by them, we rethink the assistance of segmentation applied in inter prediction, and sufficiently explore the guidance of accurate object-aware partition information for prediction and coding process in codec. In this paper, we propose an object segmentation-assisted inter prediction (SAIP) method that does not restrict the partitioning shape and further takes advantage of the segmentation information to assist the entire inter prediction process, including MC, motion vector coding (MVC), and ME.
@@ II-A2 Line-Based Geometric Partitioning Schemes
Line-based geometric partitioning schemes (GEO) have been intensively studied to capture the non-uniform motion field [10, 19, 11, 20] during the development of HEVC and VVC. In GEO[19, 11], a predefined straight line is used to split the rectangular block into two wedge-shaped sub-parts. Each sub-part has its MV to describe the motion field and performs MC to obtain the prediction. In VVC, merge-mode-based GEO with 64 partitioning modes has been adopted [11]. The mode is selected for a coding unit from the predefined 64 partition line candidates, and the description information of partition line is represented in the form of the offset parameter and angle parameter . For motion information of sub-parts, only unidirectional motion information is used, and the GEO merge list is derivated from the regular merge list by the parity of the GEO merge index. To economize the bits consumption of GEO, [20] proposed to utilize the spatio-temporal correlation to guide the efficient partitioning mode derivation and motion information coding. Although these methods [10, 19, 11, 20] can divide the motion-inconsistent sub-parts and represent their motion information with limited bits consumption in a block to perform MC, the straight line can only approximate the discontinuities of the motion field. Meanwhile, bit consumption for partition line representation is unavoidable.
@@ II-A3 Segmentation-Based Partitioning Schemes
To represent the motion field of arbitrary-shaped object accurately in natural videos, segmentation-based partitioning and coding schemes are tentatively introduced into video coding frameworks and standards [21, 22, 23, 24]. In the segmentation-based partitioning schemes, the partition is mainly derived from the reference frame/block by applying a segmentation algorithm at the encoder/decoder.
@@ III Proposed Framework
Figure 2 shows our proposed object segmentation-assisted inter prediction (SAIP) framework from the perspective of decoding. In SAIP, the key lies in that the segmentation information is utilized to assist the overall prediction and coding process in codec. It mainly consists of four modules, including MC (Section V), MVC (Section VI), rate-distortion optimization for ME and partition estimation (Section VII), segmentation for reference pictures (Section IV). The difference between the encoder and decoder is only the additional process of ME and partition estimation. In this section, we briefly introduce the process of SAIP in the decoder.
@@
In the decoding process (Fig. 2 ), different from the previous inter prediction methods, the coding block is divided into multiple motion-inconsistent regions (primary region, secondary region) to match the prediction in SAIP. First, the motion data of the current coding block is obtained by parsing the bitstream. Concerning the motion data for the coding block, two MVs (MVprimary, MVsecondary) and reference picture indexes (Refprimary, Refsecondary) correspond to the motion of different regions (primary motion, secondary motion). The partition of the coding block is translated from the segmentation mask of the reference block by the guidance of MVprimary. Second, the segmentation mask is utilized to derive the motion information of the secondary region (MVsecondary). Third, with the assistance of the segmentation mask, the final prediction can be derived by MC based on MVs of different regions.
IV Learned Segmentation for Reference Pictures
Obtaining the accurate partition of the motion area is the key to achieving efficient inter prediction. In the traditional coding process, the block-based partition is difficult to align with the irregular shape of the natural object. With the maturity of deep learning-based segmentation methods, the pixel-level demarcation representation ability described by the segmentation map is also increasing. Therefore, the combination of the segmentation method with pixel-level demarcation representation precision and the codec with block-level precision has the potential to more accurately characterize video motions along the time, thus further improving the efficiency of video coding.
Video object segmentation (VOS) aims to identify and segment target instances in a video sequence. According to the strong temporal correlation between multiple video frames, the VOS model mainly propagates the object target information of the label frame to infer the segmentation for the remaining frames. For each frame, the mask represents the segmentation information of the foreground and background. In our work, VOS is utilized to assist the partition process in obtaining the accurate partition of every object’s shape. We introduce the recent state-of-the-art segmentation method [31] [37] into codec to derive the object segmentation of the multiple reconstructed reference frames and translate the segmentation map of the reference frame to the current frame as the partition of motion-inconsistent regions. Here we focus on the details of how to embed the segmentation module into the codec, and the detailed network design follows the configurations in [31] [37].
The VOS methods can be grouped into two categories in common scenarios: unsupervised and semi-supervised. The difference is that the latter needs to obtain the annotation of the first frame as the tracking label source before the segmentation process. Due to the coding process, the segmentation process needs to be fully automatic. Although some automatic VOS methods can be automated, their segmented precision is far less accurate than that of semi-supervised VOS (SVOS). Therefore, we apply the state-of-the-art SVOS model space-time correspondence network (STCN)[37] to our framework. The STCN is a hybrid propagation-based and detection-based method based on a spatio-temporal attention mechanism, which models the complex dynamic spatio-temporal dependencies of instances in multiple frames and excellently captures the motion area to achieve accurate segmentation of instances in each frame. To replace the process of manually extracting the first frame annotations in SVOS with automatic network generation, we propose cooperation between SVOS and image instance segmentation (IIS). IIS is utilized to obtain the annotations required for the first frame of SVOS in the coding process. For IIS, we apply the state-of-the-art IIS model sample consistency network (SCNet)[31].
The derivation process of the reference picture’s segmentation map is shown in Fig. 3 . From the experiment for different segmentation configurations in the reference pictures, we find that the segmentation accuracy of IIS is influenced by the quantization parameters (QP) of the reference picture, and the VOS is not much affected by the picture quality. To fit the coding frame structure (group of pictures, GOP), SCNet is utilized to segment the first frame of the GOP as the tracking label source of SVOS and the STCN segments the other frames of the GOP.
When dealing with multi-frame segmentation in a GOP, first, SCNet is utilized to detect and segment foreground instances in the first frame of a GOP and obtain the infographic of instances, which contains the bounding box and segmentation of instances. Second, the instance segmentation information of the first frame is used as the reference for STCN inference. According to , the STCN tracks the instance in each subsequent frame and obtains of the subsequent frame. To represent different instances in the mask, the segmentation of each frame uses different colors to represent different instances and the mask of each frame is derived by binarizing . Finally, we add a buffer memory for each frame to store all to facilitate data interaction. The derivation of is expressed as follows:
where and represent the pixel values of and , and denotes the instance sets.
In the case of coding requirements, sequences are also long-term. VOS methods are especially suitable for processing short-term sequences. Although the STCN has largely solved the common challenges of VOS such as object occlusion, fast motion and deformation to a certain extent, STCN still cannot solve more complex motion very well in long-term sequences. We find many scene changes in long-term sequences with new objects appearing and camera movement. Since the STCN does not have the segmentation information of these new objects in the annotation of the first frame, it cannot segment such new objects in time, resulting in a decrease in the accuracy of the segmentation area and contour of the frame. To solve this problem, the interval of the label mask’s update can be adjusted (such as 1st-GOP, 2nd-GOP, 3rd-GOP) to adapt to the scene change of different videos. At the beginning of every segmentation interval, SCNet re-derives the object information and re-propagates the mask with the STCN, which is equivalent to updating the object information regularly to guide the accurate segmentation of SVOS.
V Segmentation-Assisted Motion Compensation
@@
As mentioned in Section IV, the reference frame can be segmented with the same setup in the encoder and decoder. The segmentation mask of the reference frame is generated when the coding frame is reconstructed, and stored in buffer memory for segmentation-assisted inter prediction of subsequent frames. With a proper indication, the object segmentation mask of the reference block is translated to the coding block as the arbitrary-shaped partition of motion-inconsistent regions. In this section, we assume that the partition and the motion information of different regions are obtained, and introduce how to utilize the segmentation mask to assist the MC for different motion-inconsistent regions (sub-regions) within the block.
@@
For different sub-regions, two MVs (MVprimary and MVsecondary) are used to represent their motion and correspond to two block-level reference blocks named primary prediction and secondary prediction (Pprimary, Psecondary). Considering the accurate representation of motion information, we propose segmentation-assisted motion compensation (SA-MC) to generate the prediction of different sub-regions within a block, the process of SA-MC is shown in Fig. 4 . In particular, the segmentation mask assists the two steps of MC, including the integer and fractional parts.
V-A Segmentation-Assisted Integer-Pixel Motion Compensation
@@
On both sides of the non-linear partition line, the 2-pixel range region is defined as the edge region. For the edge region to be predicted, not only its own MV but also the MV of the neighboring region can be used to derive the prediction signal. Fig. 5 shows a set of ORMC examples. A blending process is applied to the pixels around the non-linear partition line by a distance-based adaptive weighted strategy. Pixels are classified into three categories: one pixel away from the partition line, two pixels away, and more than two pixels away. Each category corresponds to a different weight for the fusion of Pprimary and Psecondary in every pixel.
@@
V-B Segmentation-Assisted Fractional-Pixel Motion Compensation
@@
VI Segmentation-Assisted Motion Vector Coding
@@
In SAIP, the motion of each coding block is divided into two parts (primary motion and secondary motion) to derive the motion data, which corresponds to two MVs (MVprimary, MVsecondary). One is coded by the index of the primary candidate list (length: 71), the list includes regular merge candidates (Merge) and merge with motion vector difference (MMVD) candidates [5]. The other one is coded by the index of the secondary candidate list (length: 7), the list is established with the assistance of segmentation mask.
@@
In the coding process, the primary candidate list is established at first for the coding block, and each candidate can provide a MV (MVprimary) for the current coding block’s rough motion representation. Due to multiple motions in the coding block, MVprimary can only consider describing a part of the motion. To further describe the multiple motions, the partition of motion area is needed. By the guidance of MVprimary, the segmentation mask of the primary reference block can be translated to the current coding block as the partition of the motion area. From the mask, the coding block can be partitioned into two motion-inconsistent regions (primary region, secondary region). For the motion of the primary region, it can be represented by MVprimary. In order to efficiently represent the motion of other regions, we propose a segmentation-assisted candidate derivation (SA-CD) strategy for the MV candidate derivation of other parts.
@@
The construction of the regular merge candidate list follows the fixed predefined order to traverse the potential candidates (B1A1B0A0B2T…) to fill the candidate list. Figure 6 .(i) shows the position of spatial and temporal candidates for the current coding unit (CU), which is used to copy the motion information of the neighboring block for reference. In SA-CD, the segmentation mask (object-aware partition) is utilized as the criterion to perform motion analysis for coding block and further adaptively select and rank the high-probability candidates to construct the candidate list, thereby considering different qualities of the potential MV candidates. The example is shown in Fig. 6 .(ii). First, the distribution pattern of the secondary region in the partition of the coding block’s motion-inconsistent regions is searched. The nearest four-pixel values near the corner are taken as the classification rule. In that case, the four pixels in the upper left and upper right corner of the segmentation mask are all zero (black). It is considered that these corners of the segmentation mask are in the secondary region. Similarly, the four corners can be in the secondary or the primary region, so the distribution of the motion-inconsistent regions can be classified into 16 patterns, all patterns are shown in Fig. 6 .(iii). Different patterns of distribution of motion-inconsistent regions correspond to different candidate lists. The secondary candidate list is also derived from the candidates in Fig. 6 .(i). The construction order is determined by the distribution pattern of secondary region, which assumes that the motion information of neighboring blocks closer to the secondary region is more likely to match the motion information of secondary region. Therefore, according to the distribution pattern (blue dotted box) of Fig. 6 .(ii), the order can be determined as B1B2B0(0,0)A0A1T. The candidate list of all distribution patterns is also shown in Fig. 6 .(iv).
@@
VII Segmentation-Assisted Rate-Distortion Optimization for Motion Estimation
and Partition Estimation
@@ VII Segmentation-Assisted Rate-Distortion Optimization for Motion Estimation
For the ME of different motion-inconsistent regions, the accuracy of motion representation depends on whether the partition (segmentation mask) of the motion area can approximate the actual motion field in the coding block. In the previous related work (GEO[11], PCS-16[14], TIP-19[13]), template-based or hypothesis-based partition derivation methods are utilized, the fixed search range of partition varieties and individual RDO process (for ME and partition search) limit the partition search space of coding block and the accuracy of MVs, thereby further impacting on the prediction quality. To solve these limits, segmentation-assisted rate-distortion optimization (SA-RDO) is applied to the proposed SAIP framework. In Fig. 7 , the object segmentation-assisted inter prediction framework with proposed SA-RDO (uni-directional case) is shown from the perspective of encoding. The motion candidates of MVprimary imply various partition candidates are used to suppose the motion area partition of current coding block (partition information does not need to be explicitly signaled). And the partition estimation and the ME of primary motion and secondary motion are joint to search the best combination in the whole RDO process. Here we detail the SA-RDO from the perspective of uni-directional and bi-directional prediction.
VII-A Rate-Distortion Optimization for Uni-directional Prediction